OpenStudy (anonymous):

Can someone please explain to me how to solve 0.5z+10=z+4?

OpenStudy (anonymous):

0.5z+10=z+4 remember the z on the right has a 1 in front of it, then you just subtract 0.5z from 1z to get 0.5z. 10 = 0.5z +4 Then subtract 4 from 10 which is 6. 10 = 0.5z + 4 -4 -4 6 = 0.5z Then divide 6 by 0.5 and therefore z = 12 6 = 0.5z /0.5 /0.5 z=12

OpenStudy (anonymous):

.05z+10=z+4...put all the z's on one side and all the constants on the other...so subtract a z to each side, and subtract 10 from each side...0.5z-z=-6...-0.5z=-6...divide both sides by -0.5 and z=12

OpenStudy (anonymous):

Thank Ya'll, I still don't understand how you get 0.5z when you subtracted 1z from it?

OpenStudy (anonymous):

I'm having trouble with the decimal stuff..

OpenStudy (anonymous):

its negative one half -0.5.....if its easier think of it in fractions

OpenStudy (anonymous):

Half of 1 is 0.5.

OpenStudy (anonymous):

thank ya'll!